OpenStudy (anonymous):

(bc^2)^-3 simplified and b and c are not equal...I got, 1/b^3c^6...am I correct?

OpenStudy (saifoo.khan):

\[\Huge \frac{1}{b^3c^{2 \times 3}}\]

OpenStudy (saifoo.khan):

\[\Huge \frac{1}{b^3c^{6}}\]

OpenStudy (saifoo.khan):

let me give u a medal for that.

OpenStudy (anonymous):

woohoo...sometimes I impress myself

OpenStudy (saifoo.khan):

haha.

OpenStudy (saifoo.khan):

"sometimes"